the first of two numbers is 9 more than the second. three times the first number is equal to 7 more than 5 times the second. find the numbers.
Answer by richwmiller(17219)

x=first number
y=second number
x=y+9
y=x-9
3x=7+5y
3x=7+5*(x-9)
3x=7+5x-45
3x=5x-38
38=2x
19=x
y=19-9
y=10
check
3*19=7+5*10
57=7+50
57=57